# A Grey-box Attack against Latent Diffusion Model-based Image Editing by Posterior Collapse

### v3

@@ Appendix B Gradient Direction
Notations
Throughout this section, we use the following nomenclature:
  • “PCA-” and “ours-”: Refer to minimize the $\mathcal{L}_{\text{KL}}$ (consistent with Table 1 in the main paper) with $v=1\times 10^{-8}$
  • “PCA+” and “ours+”: Refer to maximize the $\mathcal{L}_{\text{KL}}$ with $v=1$
In the subsequent tables (Table 2 and Table 3), we employ the following conventions:
  • Gray background: Indicates methods not compared with others
  • Bold text: Denotes the best results
  • Underlined text: Highlights the second-best results
Analysis of PCA+ Performance (Objective 1)
Table 2 presents a comprehensive comparison of $\operatorname{IQA}(\bm{x},f(\bm{x}_{adv}))$ across various methods, highlighting the efficacy of our PCA+ approach. The results demonstrate a consistent superior performance of our method across multiple Image Quality Assessment (IQA) metrics. Notably, PCA+ achieves the best scores in PSNR, FID, SSIM, and LPIPS across all evaluated scenarios, indicating a robust ability to maintain similarity between the edited adversarial image and the original image.
While the PCA+ method does not lead in ACDM scores, it remains highly competitive, consistently ranking among the top-tier performers. This comprehensive performance across various metrics underscores the method’s effectiveness in addressing Objective 1.
Visual analysis, as evidenced in Figures 6 and Figure 7, reveals that PCA+ produces results visually comparable to established methods such as SDS and AdvDM. A characteristic feature observed in the edited images is the appearance of watermark artifacts, which is consistent across these methods. However, PCA+ distinguishes itself through enhanced resistance to undesired edits. This superiority is not only reflected in the quantitative metrics but also apparent in qualitative analysis. A striking example is observed in the P3 scenario, where PCA+ successfully prevents the appearance of snow-like artifacts in $f(\bm{x}_{adv})$, a feat not achieved by other methods.
@@
Table 2: Results on IQAs taken between the clean image $\bm{x}$ and the adversarial sample after editing $f(\bm{x}_{adv})$. In this table, the better IQAs representing the better performance. Notbaly, “clean” for each prompt is the IQA computed between clean image $\bm{x}$ and edited clean image $f(\bm{x})$.
Analysis of PCA- Performance (Objective 2)
Table 3 provides a detailed comparison of $\operatorname{IQA}(f(\bm{x}),f(\bm{x}_{adv}))$ across different methods, showcasing the effectiveness of our PCA- approach in addressing Objective 2. The results reveal that PCA- achieves the highest level of disruption to the infringer-desired output among all evaluated methods.
Visual analysis of the results, as illustrated in Figures 6 and Figure 7, demonstrates a striking contrast between PCA- and other methods. Alternative approaches tend to preserve some degree of structural and color information similar to $f(\bm{x})$ in their $f(\bm{x}_{adv})$, however, PCA- stands out by completely obliterating the semantic information of the edited image.
A concrete example of this capability is evident in Figure 6. In this instance, where other methods produce images still recognizable as pineapples post-editing, our PCA- method generates severely distorted images. These distortions introduce semantically unrelated content, such as rubble, meaningless patterns, and even unrelated objects like trees. This level of semantic destruction effectively renders the edited image unusable for the original, potentially infringing purpose, while maintaining the integrity of the protection mechanism.
@@
P5 Ours+ 21.00 20.92 99.70 106.53 0.5402 0.5357 0.4857 0.4867 0.0548 0.0573
Table 3: Results on IQAs taken between the clean image after editing $f(\bm{x})$ and the adversarial sample after editing $f(\bm{x}_{adv})$. In this table, the worse IQAs representing the better performance.
Figure 6: A visualization with P1 = “”, P2 = “a fresh ripe pineapple is on the stem”, P3 = “add some snow”, P4 = “apply sunset lighting”, P5 = “make it like a watercolor painting”.
Figure 7: A visualization with P1 = “”, P2 = “a colorful bird is perched on a tree branch”, P3 = “add some snow”, P4 = “apply sunset lighting”, P5 = “make it like a watercolor painting”.
Appendix C Ablation Studies and Extra Evaluations
We conduct comprehensive ablation studies on all hyperparameters. It is worth noting that, all metrics taken in the ablation is following the form of PCA- (Objective 2). For the hyperparameter $v$, we also do the ablation with evaluating metrics following the form of PCA+ (Objective 1), which aims to evidence our choice of $\mathcal{N}(\mathbf{0},\mathbf{I})$ for PCA+ and $\mathcal{N}(\mathbf{0},\mathbf{0})$ for PCA-
Ablation on $v$

### v4

@@ IV-C Unified Loss Function for Two Collapse Phenomena
IV-D Optimization Framework
To craft adversarial samples, we use projected sign gradient ascent [madry2017towards] for iterative updates. Our approach is inspired by a counterintuitive finding from [sds]: they achieved better results by minimizing their loss function, which contradicts the intuitive expectation that maximization would be more effective. Following this insight, we explore both gradient directions by altering the sign of our final loss function, which will be discussed in appendix. Our update process follows the form:
$\bm{x}^{t+1}=\mathcal{P}_{\ell_{\infty}\left(\bm{\delta}^{t}\right)}^{\epsilon}\left(\bm{x}^{t}\pm\alpha\text{sign}\nabla_{\bm{x}^{t}}\mathcal{J}\left(\bm{x}^{t}\right)\right),$ (18)
where we apply plus for maximization and minus for minimization, and $\mathcal{P}_{\ell_{\infty}\left(\bm{\delta}^{t}\right)}^{\epsilon}\left(\cdot\right)$ will apply the projection on $\bm{\delta}^{t}$ in the $\epsilon$-ball of $\ell_{\infty}$ norm, $\alpha$ is a hyperparameter to adjust the learning rate during the optimization.
Algorithm 1 Posterior Collapse Attack (PCA)
0: Clean image $\bm{x}$, VAE encoder $\mathcal{E}$, perturbation budget $\epsilon$, step size $\alpha$, number of iterations $T$, variance parameter $v$, objective type $\text{obj}\in\{+,-\}$
0: Adversarial image $\bm{x}_{\text{adv}}$
1: Initialize: $\bm{x}^{0}\leftarrow\bm{x}$, $\bm{\delta}^{0}\leftarrow\mathbf{0}$
2:for $t=1$ to $T$ do
3:  // Forward pass through VAE encoder
4:  $\bm{\mu}^{t},\bm{\sigma}^{t}\leftarrow\mathcal{E}(\bm{x}^{t-1})$
5:  // Compute the loss (in Equation 12)
6:  $\mathcal{L}_{\text{KL}}(\bm{x}^{t-1})=\frac{1}{2}\sum_{i=1}^{d}\left(-\ln(\bm{\sigma}_{i}^{t})^{2}-1+\frac{(\bm{\mu}_{i}^{t})^{2}+(\bm{\sigma}_{i}^{t})^{2}}{v}\right)$
7:  // Compute gradient
8:  $\bm{g}^{t}\leftarrow\nabla_{\bm{x}^{t-1}}\mathcal{L}_{\text{KL}}(\bm{x}^{t-1})$
9:  if $\text{obj}=+$ then
10:   // Maximize for diffusion collapse (PCA+)
11:   $\bm{\delta}^{t}\leftarrow\bm{\delta}^{t-1}+\alpha\cdot\text{sign}(\bm{g}^{t})$
12:  else if $\text{obj}=$- then
13:   // Minimize for concentration collapse (PCA-)
14:   $\bm{\delta}^{t}\leftarrow\bm{\delta}^{t-1}-\alpha\cdot\text{sign}(\bm{g}^{t})$
15:  end if
16:  // Project perturbation into $\ell_{\infty}$ ball
17:  $\bm{\delta}^{t}\leftarrow\text{clip}(\bm{\delta}^{t},-\epsilon,\epsilon)$
18:  // Update adversarial image
19:  $\bm{x}^{t}\leftarrow\text{clip}(\bm{x}+\bm{\delta}^{t},0,1)$
20:end for
[…]
22:return $\bm{x}_{\text{adv}}$
Algorithm 1 presents the complete optimization procedure of our Posterior Collapse Attack. The algorithm takes a clean image $\bm{x}$ and the VAE encoder $\mathcal{E}$ as inputs, along with standard adversarial attack parameters: perturbation budget $\epsilon$, step size $\alpha$, and number of iterations $T$. Two method-specific parameters control the attack behavior: the variance parameter $v$ and the objective type $\text{obj}\in\{+,-\}$.
The optimization follows an iterative gradient-based approach. At each iteration $t$, we:
  1. 1.
    Extract the mean $\bm{\mu}^{t}$ and variance $\bm{\sigma}^{t}$ from the encoder’s output (Line 3).
  2. 2.
    Compute the loss $\mathcal{L}_{\text{KL}}$ (in Equation 12) between the posterior distribution $q(\bm{z}|\bm{x}^{t-1})$ and target distribution $p^{*}(\bm{z})=\mathcal{N}(\mathbf{0},v\mathbf{I})$ (Line 5).
  3. 3.
    Calculate the gradient with respect to the input image (Line 7).
  4. 4.
    Update the perturbation based on the objective type: gradient ascent for PCA+ ($\text{obj}=+$) to induce diffusion collapse, or gradient descent for PCA- ($\text{obj}=$-) to induce concentration collapse (Lines 8-13).
  5. 5.
    Project the perturbation into the $\ell_{\infty}$ ball to satisfy the imperceptibility constraint (Line 15).
  6. 6.
    Update the adversarial image while ensuring pixel values remain valid (Line 17).
For PCA+, we set $v=1$ to align with the standard normal prior used in VAE training, causing the posterior to diverge and create an out-of-distribution scenario. For PCA-, we set $v=1\times 10^{-8}$ to approximate zero variance, forcing the posterior to collapse to a degenerate point mass. This unified framework demonstrates the flexibility of our approach in achieving both protection objectives through a single algorithmic structure with different hyperparameter configurations.
Importantly, since our loss function $\mathcal{L}_{\text{KL}}$ operates only on the VAE encoder output ($\bm{\mu},\bm{\sigma}^{2}$), the proposed method requires no knowledge of text prompts, providing a practical advantage in real-world scenarios where the infringer’s editing intent is unknown.
IV-E Transferability Analysis
A key advantage of our method lies in its better transferability across different LDM architectures. This transferability stems from three fundamental design principles:
Minimal Model Dependency. Unlike existing methods that require white-box access to the major component of LDM, U-Net, our approach only requires access to the VAE encoder, which constitutes less than 4% of the LDM parameters. This drastically reduces the dependency on model-specific knowledge.
Architectural Commonality. The VAE encoder serves as a universal bottleneck across various LDM architectures. Most state-of-the-art LDMs, including different versions of Stable Diffusion (SD1.4, SD1.5, SD2.0, etc), share similar VAE architectures with only minor variations. By targeting this common component, our attack naturally transfers across these models.
Distribution-level Attack. Rather than attacking specific learned features or model parameters, our method manipulates the fundamental statistical properties ($\bm{\mu}$ and $\bm{\sigma}$) of the VAE’s posterior distribution. This distribution-level attack is inherently more robust to architectural variations, as any VAE-based system must process these statistical representations.
Consequently, adversarial perturbations crafted using one model’s VAE encoder (e.g., SD1.5) can effectively transfer to other LDM variants without requiring additional knowledge of the target model’s U-Net. This near-black-box setting makes our method practically applicable in real-world scenarios where full model access is unavailable.
TABLE I: Results on IQAs for Methods Subject to Objective 1.
@@ IV-E Transferability Analysis
TABLE II: Results on IQAs for Methods Subject to Objective 2.
@@
V Experiments Setup
V-A Dataset
In our experiments, we utilized a 1000-image-subset of the ImageNet [deng2009imagenet], as selected by [lin2020nesterov]. This choice aligns with established conventions in adversarial attack research. All images are $512\times 512$, ensuring consistency across evaluations.
V-B Baselines
We compared our approach against several state-of-the-art methods, AdvDM [advdm], PhotoGuard (PG) [photoguard]  [^1] We followed the choice of [sds], using the most powerful variant of PG (diffusion attack)., MIST [mist], and SDS [sds]. Those methods can be summarized into two groups: AdvDM, SDS, and PCA+ is for Objective 1; PG, MIST, and PCA- is for Objective 2. We fix fix $\alpha=2$, $T=40$ and $\epsilon=16$ for all methods, their implementations are followed with the unified framework  [^2] https://github.com/xavihart/Diff-Protect provided by SDS [sds], where the rest hyperparameters left default. We implement our method with $v=1$ for PCA+, $v=1\times 10^{-8}$ for PCA-, the optimization follows the goal defined in Equation 13 and Equation 15.
V-C Victim Models
Our experiments primarily focused on popular LDMs, specifically the lightweight Stable Diffusion 1.4 (SD14) and 1.5 (SD15). These models were chosen due to their widespread accessibility and higher potential for misuse.
To assess the transferability across different resolution and architectures, we also tested the first 100 images on Stable Diffusion 2.0 (SD20) with $768\times 768$, and Stable Diffusion XL (SDXL) with $1024\times 1024$ for all methods. All reported results of our method in transferability leverage the VAE of SD15 as the surrogate model.
V-D Varied Image Editing Prompts
To further evaluate the robustness and versatility of our method, we conducted inference using a diverse set of prompts. These prompts include:
  • P1: an empty prompt.
  • P2: a caption generated by the vision-language model [li2022blip].
  • P3: “add some snow” to represent weather modifications.
  • P4: “apply sunset lighting” to test lighting adjustments.
  • P5: “make it like a watercolor painting” to examine style transfer capabilities.
To evaluate our method’s robustness across diverse real-world editing scenarios, we additionally constructed a dataset of 200 image-prompt pairs from LAION [schuhmann2022laion] using Claude’s multimodal API333claude-sonnet-4-5-20250929. The dataset covers five editing categories: Facial Expression Change, Object Addition, Object Removal, Pose Transformation, and Background Replacement, with prompts automatically generated based on image content.
Through this varied set of prompts, we aimed to assess our method’s effectiveness across a wide range of potential image manipulation scenarios.
Figure 4: Qualitative comparison of image editing results. The first row is the original image and protected images. For the rest, each row shows different methods applied to a pineapple image under various editing prompts and protection objectives. Notably, $x$ refers to the original image will be edited by LDM; $f(\cdot)$ refers to the LDM-based image editing.
Figure 5: Qualitative comparison of image editing results. The first row is the original image and protected images. For the rest, each row shows different methods applied to a portrait photo under various editing prompts and protection objectives. Notably, $x$ refers to the original image will be edited by LDM; $f(\cdot)$ refers to the LDM-based image editing.
V-E Image Quality Assessment (IQA)
To quantitatively assess the quality of the edited images pre- and post-attack, we employed five different IQAs:
  • Peak Signal-to-Noise Ratio (PSNR): provides insight into the overall distortion introduced. [^4] Range set to 1, aligning with PyTorch’s image transformation.
  • Fréchet Inception Distance (FID): offers a measure of the perceptual quality and generation diversity [heusel2017gans]. [^5] https://pytorch.org/torcheval/stable/.
  • Structural Similarity Index (SSIM): captures changes in local patterns of pixel intensities [wang2004image]. [^6] Gaussian kernel size is 11, https://github.com/Po-Hsun-Su/pytorch-ssim.
  • Learned Perceptual Image Patch Similarity (LPIPS): measures human perceptual similarities [zhang2018unreasonable]. [^7] Encoder is VGG, https://github.com/richzhang/PerceptualSimilarity.
  • Aesthetic Color Distance Metric (ACDM): evaluates the overall color consistency [guo2024artwork, guo2026artwork]. [^8] Gaussian kernel is 11, https://github.com/ZhongliangGuo/ACDM.
V-F Evaluation Protocol
Since baselines serve different objectives, for a comprehensive performance evaluation, we employ two distinct approaches. for methods subject to Objective 1, we compute IQAs between clean images and edited adversarial samples, represented as $\operatorname{IQA}(\bm{x},f(\bm{x}_{adv}))$. for methods subject to Objective 2, we calculate IQAs between edited clean images and edited adversarial samples, denoted as $\operatorname{IQA}(f(\bm{x}),f(\bm{x}_{adv}))$.
The interpretation of IQA metrics differs between these objectives: for Objective 1, IQA metrics approaching their optimal values indicate successful protection against unauthorized editing, while for Objective 2, IQA metrics deviating from their optimal values suggest better effectiveness in disrupting unauthorized editing attempts.
V-G Against Defending Methods
To evaluate our attack’s robustness against potential defenses, we first examined Adv-Clean [advcleaner]. This defense includes multi-stage transformation including image degradations and filter blurs, which designed to remove adversarial perturbations against diffusion models. The method inherently addresses common image modifications that occur during real-world distribution and manipulation. Additionally, we evaluated these baseline methods against DiffPure, a pixel-level diffusion model-based adversarial purification method, to further assess their resilience against adversarial purification techniques. We apply these defense on our attacked images and evaluate on adversarial samples surrogated by SD15.
VI Results
Throughout this section, “PCA-” and “ours-” refer to minimize the $\mathcal{L}_{\text{KL}}$ (consistent with Tab. I) with $v=1\times 10^{-8}$; “PCA+” and “ours+” refer to maximize the $\mathcal{L}_{\text{KL}}$ with $v=1$.
In the subsequent tables, bold text denotes the best results, arrows ($\uparrow$$\downarrow$) indicate whether higher or lower values represent better protection performance. “clean” for each prompt is the IQA computed between clean image $\bm{x}$ and edited clean image $f(\bm{x})$.
VI-A Image Editing Results with Varied Prompts
TABLE III: Results on Claude-generated prompt editing for methods subject to Objective 1.
[…]
clean 15.35 153.05 0.6125 0.4077 0.0515
AdvDM 17.43 144.61 0.2719 0.6267 0.0433
SDS 17.36 146.43 0.2701 0.6300 0.0496
PCA+ 19.06 109.65 0.3910 0.6065 0.0439
TABLE IV: Results on Claude-generated prompt editing for methods subject to Objective 2.
[…]
PG 14.26 181.93 0.3973 0.6308 0.1041
MIST 14.17 194.50 0.3543 0.6213 0.0953
PCA- 13.92 166.32 0.3276 0.6468 0.1413
VI-A1 Analysis of PCA+ Performance (Objective 1)
Tab. I presents a comprehensive comparison of $\operatorname{IQA}(\bm{x},f(\bm{x}_{adv}))$ across various methods, highlighting the efficacy of our PCA+ approach.
The results demonstrate a consistent superior performance of our method across multiple Image Quality Assessment (IQA) metrics. Notably, PCA+ achieves the best scores in PSNR, FID, SSIM, and LPIPS across all evaluated scenarios, indicating a robust ability to maintain similarity between the edited adversarial image and the original image. This is further validated on the extended dataset (Tab. III) covering diverse editing categories including object removal and background replacement, where PCA+ maintains the best PSNR (19.06), FID (109.65), SSIM (0.3910), and LPIPS (0.6065).
This comprehensive performance across various metrics underscores the method’s effectiveness in addressing Objective 1.
As evidenced in Fig. 4 and Fig. 5, visual analysis reveals that PCA+ produces results visually comparable to established methods such as SDS and AdvDM. A characteristic feature observed in the edited images is the appearance of fine grid artifacts, which is consistent across these methods.
However, PCA+ distinguishes itself through enhanced resistance to undesired edits.
This superiority is not only reflected in the quantitative metrics but also apparent in qualitative analysis.
A striking example is observed in the last two rows of Fig. 4, where PCA+ successfully prevents the appearance of fox head in the dark area (penultimate row) and Graffiti-like teal artifacts (last row) in $f(\bm{x}_{adv})$, a feat not achieved by other methods.
In Fig. 5, compared to other methods, PCA+ prevent the generation of the target facial details that infringers want to make.
All visualizations showcase that the PCA+ can effectively prevent the unauthorized image editing by LDMs.
VI-A2 Analysis of PCA- Performance (Objective 2)
Tab. II provides a detailed comparison of $\operatorname{IQA}(f(\bm{x}),f(\bm{x}_{adv}))$ across different methods, showcasing the effectiveness of our PCA- approach in addressing Objective 2. The results reveal that PCA- achieves the highest level of disruption to the infringer-desired output among all evaluated methods. On the extended dataset (Tab. IV), PCA- demonstrates superior disruption with the lowest PSNR (13.92), SSIM (0.3276), and highest LPIPS (0.6468) and ACDM (0.1413) across challenging scenarios.
Visual analysis of the results, as illustrated in Fig. 4 and Fig. 5, demonstrates a striking contrast between PCA- and other methods. Alternative approaches tend to preserve some degree of structural information similar to $f(\bm{x})$ in their $f(\bm{x}_{adv})$, however, PCA- stands out by completely obliterating the semantic information of the original image.
A concrete example of this capability is evident in Fig. 4. Under the prompt “fox, Van Gogh,” other methods still maintain recognizable structural elements of the pineapple, including the outlines of both leaves and fruit portions. In contrast, PCA- almost completely deconstructs this structural information, producing images that appear more randomly generated rather than edited from the original image. Furthermore, in the last row, while other methods preserve the wall seams characteristic of normal outputs, our method not only eliminates these seams but also successfully removes the contours of the pineapple fruit section in the edited results.
Fig. 5 further demonstrates the effectiveness of PCA- across different image categories. Under various prompts, our method consistently achieves greater disruption to the generated images, pushing them further away from infringer-desired images. This is particularly evident when the prompt is “Donald Trump”, where the semantic content of the original image is barely preserved, resulting in unrecognizable outputs: a level of disruption that other methods fail to achieve.
VI-B Transferability
As illustrated in Fig. 6, our method demonstrates better transferability in achieving both protection objectives. For methods subject to Objective 1, our approach consistently achieves optimal results in PSNR, SSIM, and ACDM metrics when transferred to either SD20 or SDXL, achieving comparable performance in a black-box manner without accessing any target model parameters. Similarly, for Objective 2, our method maintains superior performance across both SD2.0 and SDXL architectures.
Notably, these performance gains are achieved with reduced model parameter access, lower runtime requirements, and decreased VRAM consumption. This suggests that our method can effectively prevent unauthorized image editing across Stable Diffusion-based models by leveraging an inherent architectural vulnerability, Posterior Collapse, in a more generalizable manner.
Figure 6: Transferability results on SD2.0 and SDXL. White-box uses the target model’s VAE; black-box uses SD1.5’s pretrained weights as surrogate. (Left) Objective 1 results. (Right) Objective 2 results. $\uparrow/\downarrow$ indicate higher/lower is better. All subplots share the same axes: x-axis shows baseline methods; y-axis shows IQA scores.
Figure 7: Analysis of KL divergence. (left) Methods subject to Objective 1, Higher values indicate stronger deviation from the standard prior. (right) Methods subject to Objective 2, Lower values indicate stronger concentration toward the origin, approaching a Dirac $\delta$ function. Values are reported as mean with std as error bar.
VI-C Analysis of Collapse Phenomena
To validate our claims about the two types of collapse in the Section IV-C, we measured KL divergence statistics for each method grouped by Objectives 1 and 2, with results shown in Fig. 7. PCA+ achieves the highest KL divergence, approximately 32% above the second-best method, empirically confirming it creates the strongest out-of-distribution effect by pushing $q(\bm{z}|\bm{x})$ furthest from $\mathcal{N}(\bm{0},\mathbf{I})$. Conversely, PCA- demonstrates the lowest KL, representing a 58% reduction compared to baselines, confirming it most effectively collapses the distribution toward a Dirac $\delta$ function.
VI-D Against Defense
As shown in Tab. V, we evaluate the robustness against Adv-Clean, which employs comprehensive low-level vision processing operations. For Objective 1, our method demonstrates superior resistance to various image degradations and filtering operations, maintaining better image preservation metrics (PSNR and SSIM) compared to baselines. For Objective 2, our approach shows stronger capability in preventing unauthorized edits, achieving significantly larger deviations from expected editing results across all metrics, particularly in FID and ACDM. This indicates our methods effectively preserve the capability to influence the image editing even under multiple image transformations including degradations and filtering blur.
The evaluation against DiffPure is shown in Tab. VI, which implements a diffusion model-based adversarial purification mechanism. For Objective 1, our approach outperforms baseline methods in preserving original image content, achieving higher scores in key image quality metrics. For Objective 2, our method maintains stronger protective capabilities compared to PhotoGuard and MIST, showing larger deviations from target editing effects. These results demonstrate that our method better preserves its protective strength against learning-based purification attempts compared to existing approaches.
This comprehensive evaluation suggests that our method maintains its effectiveness under both protection objectives even in the presence of sophisticated defense mechanisms, whether they employ traditional image processing operations or learning-based purification approaches. Notably, these superior protective capabilities are achieved with reduced computational requirements: lower VRAM consumption, faster runtime, and minimal target model parameter access, demonstrating effective protection in an almost black-box manner.
TABLE V: Defense against Adv-Clean, which evaluates the effectiveness after applying image degradation and filtering operations.
Objective 1 clean AdvDM SDS PCA+
[…]
Objective 2 PhotoGuard MIST PCA-
[…]
(Values are reported as $\text{mean}_{\pm\text{std}}$ across different prompts.)
TABLE VI: Defense against DiffPure, which evaluates the effectiveness against Diffusion-based adversarial purification.
Objective 1 clean AdvDM SDS Ours+
[…]
Objective 2 PhotoGuard MIST PCA-
[…]
(Values are reported as $\text{mean}_{\pm\text{std}}$ across different prompts.)
VII Hyperparameters Analysis
We conduct comprehensive analysis on all hyperparameters. It is worth noting that, for the hyperparameter $v$, we do the analysis with evaluation protocols subject to both Objective 1 and Objective 2, which aims to evidence our choice of $v=1$ for PCA+ and $v=10^{-8}$ for PCA-. Once the $v$ is determined, the rest analysis will follow the evaluation protocol of Objective 2 to find the optimal value.
VII-A Analysis on $v$